OpenStudy (anonymous):

Find the exact solution algebraically: 3 times 4^(x/2)=96. How would u do the work?

OpenStudy (anonymous):

first divide by 3 to get \[4^{\frac{x}{2}}=32\] and then note that \[4=2^2\] so \[4^{\frac{x}{2}}=2^x\] and then you have \[2^x=32\] which you know

OpenStudy (anonymous):

So then would you take the log of \[2^x \] and so then you would get Xlog2 and so u just divide 32 by the log of 2?

OpenStudy (mertsj):

so then you would say that 2^x = 32=2^5 so x = 5